w (4v + w) dv - 2 (v^2 - w) dw = 0


1
Expert's answer
2021-06-22T17:57:00-0400
"P(v, w)=4vw+w^2"

"P_w=4v+2w"


"Q(v,w)=-2v^2+2w"

"Q_v=-4v"

"P_w-Q_v=4v+2w+4v=8v+2w"

"\dfrac{P_w-Q_v}{P}=\dfrac{8v+2w}{4vw+w^2}=\dfrac{2}{w}=-\psi(w)"


"\mu(w)=e^{\int-{2 \over w}dw}=w^{-2}"


"(4\dfrac{v}{w}+1)dv+(-2\dfrac{v^2}{w^2}+2\dfrac{1}{w})dw=0"

"M_w=-4\dfrac{v}{w^2}"


"N_v=-4\dfrac{v}{w^2}"

"M_w=-4\dfrac{v}{w^2}=N_v"

"u(v,w)=\int(4\dfrac{v}{w}+1)dv+\varphi(w)"

"=2\dfrac{v^2}{w}+v+\varphi(w)"

"u_w=-2\dfrac{v^2}{w^2}+\varphi'(w)"

"=-2\dfrac{v^2}{w^2}+2\dfrac{1}{w}"

"\varphi'(w)=2\dfrac{1}{w}"

"\varphi(w)=2\ln w+C_1"

The general solution of the differential equation


"w (4v + w) dv - 2 (v^2 - w) dw=0"

is given by


"2\dfrac{v^2}{w}+v+\ln w^2=C"